OpenStudy (foxythepirate0984):

Is the ordered pair (1, 4) a solution to the equation y = 2x + 1? If yes, is it the only solution? A. Yes, but this ordered pair is not the only solution to the equation. B. Yes, and it is the only ordered pair that fits the equation. C. No, because the ordered pair does not fit the equation. i think B or C @autumnRoseT

OpenStudy (anonymous):

well plug 1 for x and 4 for y, what is the answer?

OpenStudy (anonymous):

is it equal?

OpenStudy (mrnood):

ok - in that pair y=4 & x =1 does y = 2 *x +1?

OpenStudy (foxythepirate0984):

4=2^1+1

OpenStudy (mrnood):

so oyu thin b or c i.e yes it is the only fit or no it doesn't fit hint - you are close!

OpenStudy (foxythepirate0984):

a..?

OpenStudy (anonymous):

so it is equal, but there are also other answer it could be, try plugging in 2 for x and 5 for y... its also equal so there are many solutions, actually it goes on for a long time

OpenStudy (anonymous):

yes A :)

OpenStudy (mrnood):

whoooa not correct

OpenStudy (foxythepirate0984):

...?

OpenStudy (mrnood):

what is th ecorect equation? \[y = 2x+1\] or \[y = 2^{x} +1\] or \[y = 2^{x+1}\]

OpenStudy (anonymous):

its the first one right @Foxythepirate0984

OpenStudy (anonymous):

I don't think she's in high enough math for the other two...

OpenStudy (mrnood):

if so then it is evidently NOT equal

OpenStudy (anonymous):

she was using this sign ^ for multiplication not as a power

OpenStudy (anonymous):

@Foxythepirate0984 I gave you the wrong answer!! they aren't equal so the answer is C not A!!

OpenStudy (anonymous):

4 isn't equal to 2(1) + 1